# On Unified Prompt Tuning for Request Quality Assurance in Public Code Review

### v1

@@ Abstract
Public Code Review (PCR) can be implemented through a Software Question Answering (SQA) community, which facilitates high knowledge dissemination. Current methods mainly focus on the reviewer’s perspective, including finding a capable reviewer, predicting comment quality, and recommending/generating review comments. Our intuition is that satisfying review necessity requests can increase their visibility, which in turn is a prerequisite for better review responses. To this end, we propose a unified framework called UniPCR to complete developer-based request quality assurance (i.e., predicting request necessity and recommending tags subtask) under a Masked Language Model (MLM). Specifically, we reformulate both subtasks via 1) text prompt tuning, which converts two subtasks into MLM by constructing prompt templates using hard prompt; 2) code prefix tuning, which optimizes a small segment of generated continuous vectors as the prefix of the code representation using soft prompt. Experimental results on the Public Code Review dataset for the time span 2011-2022 demonstrate that our UniPCR framework adapts to the two subtasks and outperforms comparable accuracy-based results with state-of-the-art methods for request quality assurance. These conclusions highlight the effectiveness of our unified framework from the developer’s perspective in public code review.
@@
Request quality assurance, Public code review, Prompt tuning
@@ 1 Introduction
2) Experimental results on the Public Code Review dataset for the time span 2011-2022 demonstrate that UniPCR outperforms comparable accuracy-based results with state-of-the-art methods for the request quality assurance.
@@ 2.1 Public Code Review Process
In this section, we outline how these request quality assurance subtasks (request necessity prediction and tag recommendation subtask) serve the public code review process [24]. The process is comprised of five core steps (as depicted in Figure 1), starting with the code writer submitting a request for change or optimization in Step 1⃝. In Step 2⃝, practitioners in the community browse and evaluate the request to assign a necessity feedback. Compliance with community norms is essential for requests to receive a response, and unnecessary requests may hinder visibility. Subsequently, in Step 3⃝, developers are asked to select several relevant technical terms to accompany their request. These tags are then used in Step 4⃝ to match the request with appropriate reviewers, who provide their feedback on the code in Step 5⃝.
Our task logic involves integrating the necessity prediction with Steps 2⃝ to provide practitioners with necessity feedback in advance. This can guide them in optimizing their request expression and improving request visibility while reducing the time spent by reviewers in Step 5⃝. Additionally, the tag recommendation is associated with Steps 4⃝, allowing requests to be correctly matched with appropriate reviewers based on the use of tags within the PCR community.
Figure 1: An Example of Public Code Review Process
@@ 3.2 Unified Prompt Tuning in PCR
In Step 1, we first parse a code segment $c$ into a data flow graph to represent the dependency relationship $Graph_{c}$ between code variables (shown on the right of Figure 2).
@@ 4.1 Dataset and Metrics
To evaluate the effectiveness of UniPCR utilizing prompt learning for the PCR (public Code Review) task, we conducted an empirical study using the dataset available which sourced from StackExchange [^1] https://archive.org/details/stackexchange, the largest public code review website, and it covers the raw data version from 2011 to 2022 [6]. In an entire PCR dataset a tag is treated as rare if its occurrence is below a predefined threshold $\theta$. Intuitively, if a tag rarely appears in such a large corpus of data, it is likely to have been a miscreated tag that is not widely recognized by developers. Therefore, these rare tags are removed, as they may reduce the quality of the training and testing data. In line with previous research [5, 4, 11], we set the threshold $\theta$ for rare tags to 50, while all requests and tags that are rare or contain only rare tags are deleted. In the end, the whole dataset contains a total of 424 tags. After cleaning, we obtained a dataset consisting of 76,161 requests. Following previous work [4, 11], the dataset is split into training, validation, and test sets using an 8:1:1 ratio.
Table 1: Statistics for public Code Review dataset
@@ 4.2 Implementation Details
We reproduce the open source code provided by authors, re-optimize the parameters for each benchmark on our dataset, and select the best result as the benchmark performance. Since comment services in PCR are often implemented as tools, our reproduction is based on their published papers.
@@ 5.3 Ablation Study
Models Accuracy F1 Unnecessary F1 Necessary
Our UniPCR 0.798 0.817 0.775
W/O Code Prefix 0.745 0.771 0.712
W/O Code & Text Prompt 0.622 0.643 0.595
@@
@3 @5 @10
Our UniPCR 0.610 0.436 0.245
w/o Code Prefix 0.565 0.404 0.228
w/o Code & Text Prompt 0.498 0.354 0.208
@@
@3 @5 @10
Our UniPCR 0.688 0.792 0.870
w/o Code Prefix 0.671 0.772 0.853
w/o Code & Text Prompt 0.586 0.673 0.773
@@
@3 @5 @10
Our UniPCR 0.619 0.541 0.372
w/o Code Prefix 0.583 0.510 0.351
w/o Code & Text Prompt 0.513 0.446 0.319
@@ 7 Conclusions and Future Work
This work proposes the UniPCR framework to jointly consider both developers and reviewers in the PCR service. We consider the request necessity prediction and tag recommendation subtasks in public code review process, and demonstrate how these two tasks serve the process. Experimental results show that our unified prompt tuning significantly outperforms state-of-the-art methods in both subtasks. Additionally, we conducted a series of ablation studies to investigate the effectiveness of text and code prompts. Finally, the UniPCR framework is promising to be transferable to other public code review tasks such as comment request necessity prediction task.
In the next step, since the UniPCR framework can also be adapted to Large Language Models (LLM), its performance on LLMs will be a promising direction for further exploration.
@@ Acknowledgments
The research is supported by the Guangxi Science and Technology Major Program (Guangxi New Energy Vehicle Laboratory Special Project: AA23062066), and National Natural Science Foundation of China: No. 62276196 and 62172311.

### v2

@@ Abstract
Public Code Review (PCR) can be implemented through a Software Question Answering (SQA) community, which facilitates high knowledge dissemination. Current methods mainly focus on the reviewer’s perspective, including finding a capable reviewer, predicting comment quality, and recommending/generating review comments. Our intuition is that satisfying review necessity requests can increase their visibility, which in turn is a prerequisite for better review responses. To this end, we propose a unified framework called UniPCR to complete developer-based request quality assurance (i.e., predicting request necessity and recommending tags subtask) under a Masked Language Model (MLM). Specifically, we reformulate both subtasks via 1) text prompt tuning, which converts two subtasks into MLM by constructing prompt templates using hard prompt; 2) code prefix tuning, which optimizes a small segment of generated continuous vectors as the prefix of the code representation using soft prompt. Experimental results on the Public Code Review dataset for the time span 2011-2023 demonstrate that our UniPCR framework adapts to the two subtasks and outperforms comparable accuracy-based results with state-of-the-art methods for request quality assurance. These conclusions highlight the effectiveness of our unified framework from the developer’s perspective in public code review.
@@
Request quality assurance, Public code review, Prompt
@@ 1 Introduction
2) Experimental results on the Public Code Review dataset for the time span 2011-2023 demonstrate that UniPCR outperforms comparable accuracy-based results with state-of-the-art methods for the request quality assurance.
@@ 2.1 Public Code Review Process
In this section, we outline how these request quality assurance subtasks (request necessity prediction and tag recommendation subtask) serve the public code review process [24]. The process is comprised of five core steps (as depicted in Figure 1), starting with the code writer submitting a request for change or optimization in Step ①. In Step ②, practitioners in the community browse and evaluate the request to assign a necessity feedback. Compliance with community norms is essential for requests to receive a response, and unnecessary requests may hinder visibility. Subsequently, in Step ③, developers are asked to select several relevant technical terms to accompany their request. These tags are then used in Step ④ to match the request with appropriate reviewers, who provide their feedback on the code in Step ⑤.
Our task logic involves integrating the necessity prediction with Steps ② to provide practitioners with necessity feedback in advance. This can guide them in optimizing their request expression and improving request visibility while reducing the time spent by reviewers in Step ⑤. Additionally, the tag recommendation is associated with Steps ④, allowing requests to be correctly matched with appropriate reviewers based on the use of tags within the PCR community.
Figure 1: An Example of Public Code Review Process (with a data sample in the middle
@@ 3.2 Unified Prompt Tuning in PCR
In Step 1, we first parse a code segment $c$ into a data flow graph to represent the dependency relationship $Graph_{c}$ between code variables (shown on the red part one the right side of Figure 2).
@@ 4.1 Dataset and Metrics
To evaluate the effectiveness of UniPCR utilizing prompt learning for the PCR (public Code Review) task, we conducted an empirical study using the dataset available which is sourced from StackExchange [^1] https://archive.org/details/stackexchange, the largest public code review website, and it covers the raw data version from 2011 to 2023 [6]. In an entire PCR dataset a tag is treated as rare if its occurrence is below a predefined threshold $\theta$. Intuitively, if a tag rarely appears in such a large corpus of data, it is likely to have been a miscreated tag that is not widely recognized by developers. Therefore, these rare tags are removed, as they may reduce the quality of the training and testing data. In line with previous research [5, 4, 11], we set the threshold $\theta$ for rare tags to 50, while all requests and tags that are rare or contain only rare tags are deleted. In the end, the whole dataset contains a total of 424 tags. After cleaning, we obtained a dataset consisting of 76,161 requests. Following previous work [4, 11], the dataset is split into training, validation, and test sets using an 8:1:1 ratio.
Table 1: Statistics for Public Code Review Dataset (Span 2011-2023)
@@ 4.2 Implementation Details
We reproduce the open source code provided by the authors, re-optimize the parameters for each benchmark on our dataset, and select the best result as the benchmark performance. Since comment services in PCR are often implemented as tools, our replication is based on their published papers.
@@ 5.3 Ablation Study
Models Accuracy F1 Unnecessary F1 Necessary $\downarrow$
Our UniPCR 0.798 0.817 0.775 -
W/O Code Prefix 0.745 0.771 0.712 5.6-8.1%
W/O Code & Text Prompt 0.622 0.643 0.595 21.2-23.2%
@@
@3 @5 @10 $\downarrow$
Our UniPCR 0.610 0.436 0.245 -
w/o Code Prefix 0.565 0.404 0.228 6.9-7.3%
w/o Code & Text Prompt 0.498 0.354 0.208 15.1-18.8%
@@
@3 @5 @10 $\downarrow$
Our UniPCR 0.688 0.792 0.870 -
w/o Code Prefix 0.671 0.772 0.853 2.0-2.5%
w/o Code & Text Prompt 0.586 0.673 0.773 11.1-15.0%
@@
@3 @5 @10 $\downarrow$
Our UniPCR 0.619 0.541 0.372 -
w/o Code Prefix 0.583 0.510 0.351 5.6-5.8%
w/o Code & Text Prompt 0.513 0.446 0.319 14.2-17.6%
@@ 7 Conclusions and Future Work
This work proposes the UniPCR framework to jointly consider both developers and reviewers in the PCR service. We consider the request necessity prediction and tag recommendation subtasks in public code review process, and demonstrate how these two tasks serve the process. Experimental results show that our unified prompt tuning significantly outperforms state-of-the-art methods in both subtasks. Additionally, we conducted a series of ablation studies to investigate the effectiveness of text and code prompts. Finally, the UniPCR framework is promising to be transferable to other public code review tasks such as comment request necessity prediction task. In the next step, since the UniPCR framework can also be adapted to Large Language Models (LLM), its performance on LLMs will be a promising direction for further exploration.
@@ Acknowledgments
The research is supported by National Natural Science Foundation of China: No. 62276196, 52031009 and 62172311, and the Guangxi Science and Technology Major Program (Guangxi New Energy Vehicle Laboratory Special Project: AA23062066).